# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ 5. SPECIAL Primitives
We now outline key secure database operations offered by SPECIAL. A major challenge in designing synopsis-assisted secure processing is the evident knowledge gap on how private synopses could potentially accelerate oblivious query processing. To bridge this gap, we explore various uses of synopses, such as creating private indexes (SPEidx 5.2) and designing compacted oblivious operations (SPEop 5.3). Moreover, as discussed earlier, joins are the most intensive operations. Hence, we extensively optimize join algorithms, synergizing both private indexing and compaction techniques to develop a novel, parallel-friendly oblivious join (§ 5.3). Another challenge in primitive design arises from ensuring lossless guarantees. To address this, we integrate mechanisms into our designs that pessimistically estimate selection cardinalities, indexing ranges, and complex join sizes, using a combination of available synopses and advanced upper bound techniques (Hertzschuch et al., 2021). For clarity, we’ll assume, WLOG., that all input relations below are of size $n$ and all 1d histograms contain $m$ bins.
@@ 5.1. Basic Operations
SPECIAL supports conventional fully-oblivious operators (Bater et al., 2017). In general, these operators function logically the same as non-private ones, but their execution transcripts are completely data independent, with results consistently padded to the worst-case maximum.
Default data access SeqACC. By default, all query executions proceed with a standard data access, loading all data into a secure array using a sequential scan [^2] In the context of a secret-shared array $\langle D\rangle$, this step can be each server sequentially load the respective secret shares of $D$ into memory, readying them for MPC.. Additionally, each loaded tuple gets a secret bit (initially ‘0’), $\mathsf{ret}$, used to mark tuple validity.
SELECT. is the secure version of filter $\sigma_{p}(R)$. It conducts a linear scan over secure array $\langle R\rangle$ for relation $R$ and updates the $\mathsf{ret}$ bit: ‘1’ for tuples satisfying predicate $p$ and ‘0’ for others. Every tuple’s $\mathsf{ret}$ bit is updated, regardless of whether it matches the predicate.
PROJECT. discards irrelevant attributes for relation stored in $\langle R\rangle$. Note that the secret bit $\mathsf{ret}$ cannot be discarded by a PROJECT.
JOIN. implements the secure version of $\theta$-join, $R_{0}\Join_{\theta}R_{1}$. It processes two secure arrays $\langle R_{0}\rangle$ and $\langle R_{1}\rangle$, computing their cartesian product $\langle R_{0}\times R_{1}\rangle$ and then using SELECT to mark the joined tuples. It’s important to note that JOIN expands output, always padding it to the worst-case max (product of input sizes).
COUNT, SUM, MIN/MAX are the aggregation primitives. The operators linearly scan a secure array and continuously update a secret-shared aggregation value after each tuple access.
ORDER-BY, DISTINCT, GROUP-BY(AGG). are built on the oblivious sort primitive (Batcher, 1968). Specifically, ORDER-BY obliviously sorts a secure array according to a given attribute. DISTINCT sorts a secure array first, then performs a linear scan to identify unique tuples. For instance, among consecutively sorted identical tuples, it sets the $\mathsf{ret}$ bit of the last tuple to ‘1’ and the rest to ‘0’. GROUP-BY(AGG) starts by using DISTINCT to identify unique tuples. For each distinct tuple ($\mathsf{ret}$ set to ‘1’), it appends an aggregation value derived from the tuple and a dummy attribute (e.g., ‘-1’) for non-distinct tuples.
@@ 5.2. SPECIAL Index SPEidx
In conventional databases, indexes are powerful data structures that map attribute values to positions within a sorted array, enabling fast data access. However, generalizing such feature to SCA remains largely underexplored with existing solutions facing significant limitations such as high data loss (Roy Chowdhury et al., 2020), dependence on complex data structures filled with dummy tuples (Zhang et al., 2023; Bogatov et al., 2021), and limited query capabilities (i.e., linear queries across fixed attributes). Moreover, these methods only index base relations. SPEidx innovatively overcomes these issues by enabling the creation of lossless indexes directly on outsourced data or query intermediate results, without the need for additional structures or dummy injections.
In general, SPEidx builds upon the typical indexing model that utilizes cumulative frequencies (CF) (Kraska et al., 2018). Specifically, given $D$ sorted by $A\in attr(D)$, all records $t\in D$ where $t.A=x$ can be indexed by the interval $[g(x-1),g(x)]$, where $g(x)=|\{t\mid t.A\leq x\}|$ is the CF function. For better illustration, we show an example index lookup in Figure 2: to get all records with an attribute value of 24, one may compute $[g(23),g(24)]=[217,248]$ and access the relevant data from the subset $D[217:248]$. To make this indexing method private and lossless, the key idea of SPEidx is to derive two noisy CF curves from synopses (e.g. bounding histograms). One curve, $g^{+}(x)$, consistently overestimates $g(x)$, while the other, $g^{-}(x)$, consistently underestimates it. Then for any attribute value $x$, we can now derive a private interval $[g^{-}(x-1),g^{+}(x)]$ that losslessly indexes all desired records. For instance, as illustrated in Figure 2, the SPECIAL index might estimate the index range for attribute value 24 as $[g^{-}(23),g^{+}(24)]=[198,267]$.
Figure 2. True index (left) vs. SPECIAL index (right) for $x=24$
@@
Given $D$ sorted by $A$, the bounding histogram $\mathbf{H}(A,D)=\{\mathbf{h}^{+},\mathbf{h}^{-}\}$, and assume $\mathbf{h}^{+}=(c^{+}_{1},...,c^{+}_{m})$, $\mathbf{h}^{-}=(c^{-}_{1},...,c^{-}_{m})$ partitions $dom(A)$ into $\{B_{1},...,B_{m}\}$. We say $\textsc{SPEidx}(A,D)=\{\mathsf{idx}_{i}=[\mathsf{lo}_{i},\mathsf{hi}_{i}]\}_{1\leq i\leq m}$ is the SPECIAL index of $D$ over $A$ with:
  • $\forall~{}i\geq 1$, $\mathsf{hi}_{i}=\min(|D|,\sum^{i}_{k=1}c^{+}_{k})$.
@@
By this construction, all tuples $t\in D$ such that $t.A\in B_{i}$ will be organized into the subset $D[\mathsf{idx}_{i}]\subseteq D$. This subset can be quickly accessed if $D$ is already sorted, without the need for special data structures or inclusion of dummy tuples. Depending on how bounding histograms are constructed, $\textsc{SPEidx}(A,D)$ can support indexing lookups with varying granularity. This can range from indexing individual attribute values (where each $B_{i}$ corresponds to a single domain value) to indexing a range of of values. The bounding histogram’s pessimistic estimation ensures that all tuples where $t.A\in B_{i}$ are accurately contained within $D[\mathsf{idx}_{i}]$, thereby achieving lossless indexing. In contrast to existing methods that are limited to indexing base relations (Roy Chowdhury et al., 2020; Bogatov et al., 2021; Zhang et al., 2023), SPEidx extends its capabilities to create private indexes on query intermediate results. For instance, consider $D^{\prime}\leftarrow\sigma_{A^{*}\in vals}(D)$ where the attribute pair $(A^{*},A)$ is included in $\mathtt{synop}$. Here, SPEidx can derive $\mathbf{H}(A,D^{\prime})$ from $\mathbf{H}(A^{*},D)$ and subsequently build indexes on $D^{\prime}$. Importantly, since index creation is a post-processing activity using available DP synopses, it incurs no additional privacy loss.
Indexed store and fast data access IdxAcc. SPEidx enables the exploration of a new storage layout for organizing outsourced data, namely indexed datastore, that facilities private data access. Specifically, by analyzing a representative workload $Q_{\mathsf{R}}$, one may identify the “hottest” attribute per base relation, sort them according to the “hottest” attribute, and then build indexes over the sorted data. This storage layout enables fast indexed access (IdxAcc) to retrieve a compact subset of data from the outsourced relations, thereby eliminating the need for a full table sequential scan (SeqAcc) potentially bypassing selection operations (§ 5.3). We emphasize that the SPECIAL design does not require replicating the outsourced datastore to accommodate multiple query types (Bogatov et al., 2021; Zhang et al., 2023). However, creating compact replicas (e.g., column replicas (Huang et al., 2020) over frequently queried attributes) can be optionally employed to enhance query processing speed. Moreover, the generation of all aforementioned objects (indexed store and column replicas) requires only three primitives: projection, oblivious sorting, and SPEidx, all achieved without extra privacy loss. In other words, this implies that one can selectively adjust these objects to align with dynamic query workloads, without incurring any privacy costs.
@@ 5.3. SPECIAL Operators SPEop
We introduce SPEop, a set of novel synopsis-assisted operators that maintain full obliviousness (e.g., operators’ execution cause no privacy loss) while enabling lossless compaction. To our knowledge, SPEop is the first primitive of its kind in any SCA system.
Oblivious compaction: OPAC. We first introduce the fundamental operation of oblivious compaction, a key component for many SPEop primitives. Given input $\langle R\rangle$, OPAC sorts it based on the secret bit $\mathsf{ret}$, moving tuples with $\mathsf{ret}=$ ‘1’ to the front. Then, OPAC retains only the first $k$ tuples from the sorted array. The compaction is lossless if $k$ is greater than or equal to the number of tuples with $\mathsf{ret}=$ ‘1’; otherwise, it is lossy.
@@
(SP)SELECT. The running complexity of (OP)SELECT depends on OPAC, which is typically linearithmic (see § 6.1 or (Sasy et al., 2022)). However, when $\mathtt{CardEst}(\sigma_{A\in vals}(R),\mathtt{synop})$ is relatively small, oblivious selection can be achieved without necessarily incurring linearithmic cost. Specifically, we consider (SP)SELECT, which first creates an empty output array $\langle R_{o}\rangle$ with size equals to $cs$ before any computations. Next, it evaluates two linear scans over $\langle R\rangle$, where the first scan obliviously marks all selected tuples, and in the second scan, it privately writes all marked tuples into $\langle R_{o}\rangle$. Specifically, in the second scan, (SP)SELECT internally maintains the last actual write position $\mathsf{idx}$ in $\langle R_{o}\rangle$. Then for every newly accessed tuple $\langle t\rangle$ in $\langle D\rangle$, a write action occurs on all tuples in $\langle R_{o}\rangle$. If $\langle t\rangle$ is selected, then an actual write is made that writes $\langle t\rangle$ to $\langle R_{o}[\mathsf{idx}+1]\rangle$ and a dummy write [^3] In the context of the secret-shared secure array $\langle\mathbf{a}\rangle$, a dummy write to $\langle\mathbf{a}[i]\rangle$ is simply a re-sharing of $\mathbf{a}[i]$ through secure protocols without changing its value. is made to elsewhere. If not, dummy writes are made throughout $\langle R_{o}\rangle$. Similarly, (SP)SELECT does not yield additional privacy loss, and the complexity is linear in its input.
(DC)SELECT. Finally, if the underlying data is already indexable on $A$, a direct selection over the data can be applied that bypasses secure computations. The operator simply looks up $\textsc{SPEidx}(A,R)$, and accesses $R[a,b]$, where $a=\min_{i}(\mathsf{idx}_{i}.\mathsf{lo}),b=\max_{i}(\mathsf{idx}_{i}.\mathsf{hi})$, where $\mathsf{idx}_{i}$ dentoes the index in $\textsc{SPEidx}(A,R)$ where $B_{i}\cap vals\neq\emptyset$.
SPECIAL join: (MX)JOIN. We now introduce a novel MF-Index based oblivious join operation. The advancements of (MX)JOIN stand out in two key aspects compared to existing oblivious joins. First, compared to the standard JOIN, (MX)JOIN stands out for its ability to significantly compact the output size, coupled with a highly parallelizable fast processing mode. Second, existing oblivious joins with DP leakages typically require additional privacy budget to learn join sensitivity (Dong et al., 2022) or necessitate truncations on joined tuples (Wang et al., 2022; Bater et al., 2018). Moreover, (MX)JOIN is unique as the first oblivious join that enables lossless output compaction without extra privacy loss. We illustrate the construction details in Algorithm 3.
@@
1:Input: relations $R_{0}$, $R_{1}$; join attribute $A_{\mathsf{j}}$.
@@
10:$\mathsf{BucketJoin}(R_{0},R_{1},A_{\mathsf{j}})$:
11:for $i=1,2,...,m$ do
12:     compute $O_{i}\leftarrow(R_{0}[\mathsf{idx}_{i}]\Join_{A_{\mathsf{j}}}R_{1}[\mathsf{idx}_{i}])$ via basic JOIN
13:     $cs_{i}\leftarrow\min\left(\frac{|R_{0}[\mathsf{idx}_{i}]|}{\widehat{\mathtt{mf}}(A_{\mathsf{j}},R_{0})},\frac{|R_{1}[\mathsf{idx}_{i}]|}{\widehat{\mathtt{mf}}(A_{\mathsf{j}},R_{1})}\right)\times{\widehat{\mathtt{mf}}(A_{\mathsf{j}},R_{0})}\cdot{\widehat{\mathtt{mf}}(A_{\mathsf{j}},R_{1})}$
14:     $R_{\mathsf{out}}\leftarrow R_{\mathsf{out}}\cup{\texttt{OPAC}}(O_{i},cs_{i})$
15:return $R_{\mathsf{out}}$
In general, (MX)JOIN can be applied to two types of data: the base and pre-filtered relations where the join key attribute is contained by the synopses. Specifically, (MX)JOIN starts with computing the join key MFs (Alg 3:4) and constructing private indexes (Alg 3:5) for both inputs. All these operations are conducted through “privacy cost-free” transformations using available DP synopses. Once these objects are obtained, the algorithm employs oblivious sort to rearrange both inputs (Alg 3:6,7), rendering them indexable with tuples logically distributed into independent buckets by join key values. Next, (MX)JOIN simply adopts standard JOIN to join tuples exclusively within the same buckets (Alg 3:10). Finally, (MX)JOIN performs per-bucket output compaction, where it first determines the MF join bound (Hertzschuch et al., 2021) for each bucket join and invokes OPAC to compact the output according to the learned size (Alg 3:11,12). As bucket-wise operations are independent, the aforementioned steps lend themselves well to parallelized processing. For better illustration, we visualize (MX)JOIN’s processing flow in Figure 3.
Figure 3. Example processing flow of (MX)JOIN.
Since (MX)JOIN derives join compaction sizes completely from post-processing of DP synopses, it thus incurs no extra privacy loss. Moreover this design choice also eliminates the need for computing join sensitivities — a major hurdle in existing — as accurate join sensitivity estimation is hard and typically requires a significant privacy budget (Dong et al., 2022). Additionally, the noisy MF bounds guarantee that compaction sizes are consistently overestimated, ensuring lossless compaction of join results.

### v2

@@ 5. SPECIAL Primitives
We next introduce the secure primitives in SPECIAL. One major challenge in designing these primitives is the knowledge gap on how private synopses can accelerate oblivious query processing, i.e., C-3. To address this, we explore various usage of synopses, including creating private indexes (SPEidx § 5.2) and designing compacted oblivious operations (SPEop § 5.3). Given that joins are the most resource-intensive operations, we optimize join algorithms by combining private indexing and compaction techniques to develop a novel, parallel-friendly oblivious join (§ 5.3). Another challenge is ensuring lossless processing, which we tackle by integrating mechanisms that pessimistically estimate selection cardinalities, indexing ranges, and join sizes using synopses (§ 4) and advanced upper bound techniques (Hertzschuch et al., 2021). For simplicity, we assume all input relations are of size $n$ and all 1D histograms contain $m$ bins.
@@ 5.1. Basic Operations
SPECIAL supports conventional fully-oblivious operators (Bater et al., 2017), which logically the same as non-private ones but with data-independent execution and worst-case padding for results. We briefly introduce these operations: (i) Default data access (SeqACC). By default, query execution begins with loading all data into a secure array via sequential scan. Each loaded tuple gets a secret bit $\mathsf{ret}$ (initially ‘0’), marking its validity; (ii) SELECT. This secure filter $\sigma_{p}(R)$ performs a linear scan over the secure array $\langle R\rangle$, updating the $\mathsf{ret}$ bit to ‘1’ for tuples satisfying predicate $p$ and ‘0’ for others; (iii) PROJECT. Removes irrelevant attributes from relation $\langle R\rangle$, but retains the $\mathsf{ret}$ bit; (iv) JOIN. Implements a secure $\theta$-join $R_{0}\Join_{\theta}R_{1}$ by computing the cartesian product $\langle R_{0}\times R_{1}\rangle$ and marking joined tuples with SELECT. The output is padded to the worst-case maximum size; (v) COUNT, SUM, MIN/MAX. These aggregation operators scan the secure array and update a secret-shared aggregation value for each tuple; (vi) ORDER-BY, DISTINCT, GROUP-BY(AGG). Built on the oblivious sort primitive (Batcher, 1968). ORDER-BY sorts the array by a given attribute. DISTINCT sorts and then identifies unique tuples, marking only the last in a sequence of identical tuples with $\mathsf{ret}=\text{`1'}$. GROUP-BY(AGG) first uses DISTINCT to find unique tuples. For each distinct tuple ($\mathsf{ret}$ set to ‘1’), it appends an aggregation value derived from the tuple and a dummy attribute (e.g., ‘-1’) for non-distinct tuples.
@@ 5.2. SPECIAL Index SPEidx
Existing index techniques in SCA have several drawbacks: loss of qualified data (Roy Chowdhury et al., 2020), reliance on intricate data structures with large overhead (Zhang et al., 2023; Bogatov et al., 2021; Qiu et al., 2023), and restricted query support (Zhang et al., 2023; Bogatov et al., 2021). Furthermore, all these techniques only support indexing the base relations. SPEidx offers a breakthrough by enabling the creation of lossless indexes directly on outsourced data and the query of intermediate results, eliminating the need for extra structures or storing dummy data.
In general, SPEidx builds upon the typical indexing model that utilizes cumulative frequencies (CF) (Kraska et al., 2018). Specifically, given $D$ sorted by $A\in attr(D)$, all records $t\in D$ where $t.A=x$ can be indexed by the interval [g(x1),g(x)][g(x-1),g(x)], where $g(x)=|\{t\mid t.A\leq x\}|$ is the CF function. For better illustration, we show an example index lookup in Figure 2: to get all records with an attribute value of 24, one may compute $[g(23),g(24)]=[217,248]$ and access the relevant data from the subset $D[217:248]$. To make this indexing method private and lossless, the key idea of SPEidx is to derive two noisy CF curves from synopses (i.e., bounding histograms). One curve, $g^{+}(x)$, consistently overestimates $g(x)$, while the other, $g^{-}(x)$, consistently underestimates it. Then for any attribute value $x$, we can now derive a private interval $g^{-}(x-1),g^{+}(x)]$ that losslessly indexes all the desired records. As illustrated in Figure 2, the SPECIAL index might estimate the index range for attribute value 24 as $g^{-}(23),g^{+}(24)]=[198,267]$.
Figure 2. True (left) vs. SPECIAL (right) index for $x=24$
@@
[…]
@@
By this construction, all tuples $t\in D$ such that $t.A\in B_{i}$ will be organized into the subset $D[\mathsf{idx}_{i}]\subseteq D$. This subset can be quickly accessed if $D$ is already sorted, without the need for special data structures or inclusion of dummy tuples. Depending on how bounding histograms are constructed, $\textsc{SPEidx}(A,D)$ can support indexing lookups with varying granularity. This can range from indexing individual attribute values (where each $B_{i}$ corresponds to a single domain value) to indexing a range of of values. The bounding histogram’s pessimistic estimation ensures that all tuples where $t.A\in B_{i}$ are accurately contained within $D[\mathsf{idx}_{i}]$, thereby achieving lossless indexing. In contrast to existing methods that are limited to indexing base relations (Roy Chowdhury et al., 2020; Bogatov et al., 2021; Zhang et al., 2023), SPEidx extends its capabilities to create private indexes on query intermediate results. For instance, consider $D^{\prime}\leftarrow\sigma_{A^{*}\in vals}(D)$ where the attribute pair $(A^{*},A)$ is included in $\mathtt{synop}$. Here, SPEidx can derive $\mathbf{H}(A,D^{\prime})$ from $\mathbf{H}(A^{*},D)$ and subsequently build indexes on $D^{\prime}$. Importantly, since index creation is a post-processing procedure using available DP synopses, it incurs no additional privacy loss.
Indexed store and fast data access IdxAcc. SPEidx enables a new storage layout for outsourced data, namely indexed datastore. Specifically, by analyzing a representative workload $Q_{\mathsf{R}}$, one may identify the “hottest” attribute per base relation, sort them according to the “hottest” attribute, and then build indexes over the sorted data. This storage layout enables fast indexed access (IdxAcc) to retrieve a compact subset of data from the outsourced relations, thereby eliminating the need for a full table sequential scan (SeqAcc) and can directly produce a compact input. We emphasize that the SPECIAL design does not require replicating the outsourced datastore to accommodate multiple query types (Bogatov et al., 2021; Zhang et al., 2023). However, creating compact replicas (e.g., column replicas (Huang et al., 2020) over frequently queried attributes) can be optionally employed to enhance query processing speed. Moreover, the generation of all aforementioned objects (indexed store and column replicas) requires only three primitives: projection, oblivious sorting, and SPEidx. In other words, this implies that one can selectively adjust these objects to align with dynamic query workloads, without incurring extra privacy loss.
@@ 5.3. SPECIAL Operators SPEop
We introduce SPEop, a set of novel synopsis-assisted operators that maintain full obliviousness, while enabling lossless compaction. To our knowledge, SPEop is the first primitive of its kind in any SCA.
Oblivious compaction: OPAC. is a fundamental operation critical to other SPEop primitives. Given input $\langle R\rangle$, OPAC sorts it based on the secret bit $\mathsf{ret}$, moving tuples with $\mathsf{ret}=$ ‘1’ to the front. Then, OPAC retains only the first $k$ tuples from the sorted array. The compaction is lossless if $k$ is greater than or equal to the number of tuples with $\mathsf{ret}=$ ‘1’; otherwise, it is lossy.
@@
(SP)SELECT. The running complexity of (OP)SELECT depends on OPAC, which is typically linearithmic (see § 6.1 or (Sasy et al., 2022)). However, when $\mathtt{CardEst}(\sigma_{A\in vals}(R),\mathtt{synop})$ is relatively small, oblivious selection can be achieved without necessarily incurring linearithmic cost. Specifically, we consider (SP)SELECT, which first creates an empty output array $\langle R_{o}\rangle$ with size equals to $cs$ before any computations. Next, it evaluates two linear scans over $\langle R\rangle$, where the first scan obliviously marks all selected tuples, and in the second scan, it privately writes all marked tuples into $\langle R_{o}\rangle$. Specifically, in the second scan, (SP)SELECT internally maintains the last actual write position $\mathsf{idx}$ in $\langle R_{o}\rangle$. Then for every newly accessed tuple $\langle t\rangle$ in $\langle R\rangle$, a write action occurs on all tuples in $\langle R_{o}\rangle$. If $\langle t\rangle$ is selected, then an actual write is made that writes $\langle t\rangle$ to $\langle R_{o}[\mathsf{idx}+1]\rangle$ and a dummy write is made to elsewhere. If not, dummy writes are made throughout $\langle R_{o}\rangle$. We say that, in the context of the secret-shared secure array $\langle\mathbf{a}\rangle$, a dummy write to $\langle\mathbf{a}[i]\rangle$ is simply a re-sharing of $\mathbf{a}[i]$ through secure protocols without changing its value.
(DC)SELECT. Finally, if the underlying data is already indexable on $A$, a direct pre-fetch can be applied to avoid full table scan and compaction. The operator simply looks up $\textsc{SPEidx}(A,R)$, and accesses $R[a,b]$, where $a=\min_{i}(\mathsf{idx}_{i}.\mathsf{lo}),b=\max_{i}(\mathsf{idx}_{i}.\mathsf{hi})$, and $\mathsf{idx}_{i}$ dentoes the index in $\textsc{SPEidx}(A,R)$ with bin $B_{i}\cap[a,b]\neq\emptyset$. A standard SELECT is then applied to $R[a,b]$.
SPECIAL join: (MX)JOIN. We now introduce a novel MF-Index based oblivious join operation. The advancements of (MX)JOIN stand out in three aspects. First, compared to the standard JOIN, (MX)JOIN stands out for its ability to significantly compact the output size, coupled with a highly parallelizable fast processing mode. Second, existing DP oblivious joins typically require spending privacy budget (Dong et al., 2022) to learn join sensitivity (Dong et al., 2022) or necessitate truncation on joined tuples (Wang et al., 2022; Bater et al., 2018). (MX)JOIN eliminates this need. Moreover, (MX)JOIN is unique as the first oblivious join that enables lossless output compaction without extra privacy loss. We illustrate the construction details in Algorithm 3.
@@
1:Input: relations $R_{0}$, $R_{1}$; join attribute $A_{\mathsf{j}}$; we consider synopses (histograms) of $A_{\mathsf{j}}$ are partitioned into bins $B_{1},...B_{m}$.
@@
[…]
11:for i=1,2,,mi=1,2,...,m do
12:  $R^{(i)}_{0,1}\leftarrow\sigma_{A_{j}\in B_{i}}(R_{0,1})$ using (DC)SELECT/ $\textsc{SPEidx}(A_{\mathsf{j}},R_{0,1})$
13:  compute $O_{i}\leftarrow(R^{(i)}_{0}\Join_{A_{\mathsf{j}}}R^{(i)}_{1})$ via standard JOIN
14:  $cs_{i}\leftarrow\min\left(\frac{|R^{(i)}_{0}|}{\widehat{\mathtt{mf}}(A_{\mathsf{j}},R_{0})},\frac{|R^{(i)}_{1}|}{\widehat{\mathtt{mf}}(A_{\mathsf{j}},R_{1})}\right)\times{\widehat{\mathtt{mf}}(A_{\mathsf{j}},R_{0})}\cdot{\widehat{\mathtt{mf}}(A_{\mathsf{j}},R_{1})}$
15:  $R_{\mathsf{out}}\leftarrow R_{\mathsf{out}}\cup{\texttt{OPAC}}(O_{i},cs_{i})$
16:return $R_{\mathsf{out}}$
In general, (MX)JOIN can be applied to two types of data: the base and pre-filtered relations where the join key attribute is included in $\mathtt{synop}$. Specifically, (MX)JOIN starts with computing the join key MFs (Alg 3:4) and constructing private indexes (Alg 3:5) for both inputs. All these operations are conducted through “privacy cost-free” transformations using available DP synopses. Once these objects are obtained, the algorithm employs oblivious sort to rearrange both inputs (Alg 3:6,7), rendering them indexable with tuples logically distributed into independent buckets by join key values. Next, (MX)JOIN simply adopts standard JOIN to join tuples exclusively within the same buckets (Alg 3:10). Finally, (MX)JOIN performs per-bucket output compaction, where it first determines the MF join bound (Hertzschuch et al., 2021) for each bucket join and invokes OPAC to compact the output according to the learned size (Alg 3:11,12). As bucket-wise operations are independent, the aforementioned steps lend themselves well to parallelized processing. As (MX)JOIN derives join compaction sizes completely from post-processing of DP synopses, it thus incurs no extra privacy loss. Additionally, the noisy MF bounds guarantee that compaction sizes are consistently overestimated, ensuring lossless compaction of join results.